How many grams of Kr are in a 7.46 L cylinder at 60.2 ∘C and 6.39 atm?
Answer #1

= 7.464 T= (60.2 + 273.15 lk T= 333.35k P= 6.39 atm R = 0.0821 L. atm. met! k-1 molar mass of kr M = 83.798 g.mol-1 ideal gas